Calculus II, Spring 2009 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 12

Assigned May 14, 2009.

1. Section 14.7: Problems 3, 5, 8(can you figure out how one chooses the change of variables u, v in this problem?) , 11, 13.

2. Chap 14: 33.